# Large Model Based Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

### v2

@@ III-E2 Distributed Competition
  • MAD. The cognitive behavior of LLMs, such as self-reflection, has proven effective in solving NLP tasks but can lead to thought degeneration due to biases, rigidity, and limited feedback. LM agents in MAD [104] involve engaging in structured arguments or debates to defend their positions and challenge the strategies of others. MAD explores interactions among LLMs, where agents engage in a dynamic tit-for-tat, allowing them to correct each other’s biases, overcome resistance to change, and provide mutual feedback. In MAD, diverse role prompts (i.e., personas) are crucial and there are generally three communication strategies: (a) one-by-one debate, (b) simultaneous-talk, and (c) simultaneous-talk-with-summarizer, as illustrated in Fig. 17. In one-by-one debate strategy, debater agents respond sequentially in a fixed order, generating replies based on the ongoing discussion. Each agent’s chat history is updated with previous responses to maintain context. In simultaneous-talk debates, debater agents respond asynchronously within each round, mitigating the influence of speaking order on the debate’s flow. The simultaneous-talk-with-summarizer debate strategy enhances simultaneous-talk by incorporating a summarizer LLM. After each round, the summarizer condenses the discussion into a cohesive summary, which replaces the chat histories of all agents, ensuring shared context.
  • Non-cooperative Game. In non-cooperative games, LM agents engage in strategic decision-making where each LM agent’s goal is to seek the Nash equilibrium. Various non-cooperative games can be employed such as Nash bargaining, auctions, matching games. For instance, Xu et al. [105] design a new cached model-as-a-resource paradigm and introduce the concept of age-of-thought to efficiently deploy LLM agents. They also devise a deep Q-network-based auction mechanism to incentivize network operators.
@@ III-E4 Lessons Learned
In distributed collaboration, SL and MoE focus on model segmentation, while FL and role-playing emphasize model integration. Within the overall cloud-edge collaboration for LM agent systems, key challenges emerge in the joint consideration of LM segmentation (e.g., SL and MoE), LM placement, LM replacement, LM routing, and knowledge caching, alongside the inherent heterogeneity and dynamics of edge environments. Effective deployment and operation of LM agents across cloud and edge infrastructures require a delicate balance between LM segmentation and LM integration strategies, with an emphasis on optimizing resource allocation and minimizing service latency. Additionally, solutions should be adaptable to the constantly changing network conditions and hardware limitations inherent to edge devices. Collaborative models such as SL and MoE can significantly reduce computational burdens on edge devices, while federated strategies and role-playing can enhance cooperation among distributed agents. Future research should explore the seamless integration of these techniques, taking into account both static and dynamic aspects of edge networks, to provide more efficient and scalable solutions for large-scale applications.
@@ III-F Cooperation Paradigms of Connected LM Agents
As illustrated in Fig. 18, the detailed cooperation paradigms of LM agents under cloud-edge-end architecture involves three perspectives: data cooperation, computation cooperation, and knowledge cooperation.
  • Data Cooperation: Within a common task, LM agents continuously exchange and fuse their individual data (e.g., task-oriented sensory data) to ensure a comprehensive and up-to-date understanding of their task environment, thereby enhancing collective intelligence and enable coordinated actions.
@@
a) Multimodal Data Cooperation emphasizes the fusion of data from various modalities, such as text, images, audio, and video, to provide a comprehensive understanding of the environment. This cooperation allows LM agents to process and interpret information from multiple sources, leading to more accurate and robust decision-making. (i) By combining data from different modalities, it helps to create a unified representation that takes advantage of the strengths of each type of data. For instance, Gross et al. [108] discuss the use of multimodal data to model communication in artificial social agents, emphasizing the importance of verbal and nonverbal cues for natural human-robot interaction. (ii) By enabling LM agents to retrieve relevant information across different modalities, it enhances their ability to respond to complex queries and scenarios. For instance, by considering the intra-modality similarities in multi-modal video representations, Zolfaghari et al. [109] introduce the contrastive loss in contrastive learning process for enhanced cross-modal embedding. The effectiveness is validated using LSMDC and Youcook2 datasets for video-text retrieval task and video captioning task.
@@
a) Horizontal collaboration refers to decompose complex tasks into manageable sub-tasks, with multiple LM agents independently completing their assigned tasks in parallel. The results are then summarized and integrated to generate the final outcome. This approach enables the collaborative system to scale horizontally by adding more LM agents, allowing for more complex and dynamic tasks to be handled. The parallel processing and multi-angle perspectives contribute to increased robustness and reduced errors and biases that may arise from relying on a single agent. As shown in Fig. 19(a), each LM agent independently analyzes different aspects of stock performance such as trend, news, and industry to assess its investment potential. Horizontal collaboration has been successfully implemented in various LM agent systems. For instance, ProAgent [84] employs a framework where LM agents collaborate as teammates, analyzing each other’s intentions and updating their beliefs based on observed behaviors of their peers. This enhances collective decision-making by allowing agents to adapt dynamically in real time. Similarly, DyLAN [106] assembles a team of strategic agents that communicate through task-specific queries, allowing multiple rounds of interaction to enhance both efficiency and overall performance. However, the independent analysis of each LM agent may lead to inconsistent outputs, complicating the aggregation of the final conclusions. Therefore, effective mechanisms are needed to address disagreements and ensure that the agents’ contributions are complementary, which may increase the complexity of the coordination process.
@@
b) Vertical collaboration involves decomposing complex tasks into multiple stages, with different LM agents handling each stage sequentially. After completing their respective tasks, each agent passes the result to the next agent, until the entire task is successfully completed. As illustrated in Fig. 19(b), in medical image analysis, the first agent might extract basic visual features, the second agent could analyze these features in greater depth, and the final agent produces a diagnosis. This sequential approach enables a step-by-step refinement, allowing complex problems to be broken down and tackled more efficiently by leveraging specialized agents at each stage. For instance, Jiang et al. [24] present CommLLM, a multi-agent system for natural language-based communication tasks. It comprises three components: 1) multi-agent data retrieval, which uses condensate and inference agents to refine 6G communication knowledge; 2) multi-agent collaborative planning, employing various planning agents to generate solutions from multiple viewpoints; and 3) multi-agent reflection, which evaluates solutions and suggests improvements through reflexion and refinement agents. However, a limitation of vertical collaboration is that higher-level LM agents rely on the accuracy of lower-level outputs, making the process vulnerable to error propagation. Mistakes made early in the sequence can compromise the final outcome, highlighting the need for high precision at each stage to maintain the overall system’s reliability.
c) Hybrid collaboration. In practical LM agent environments, real-world applications often require a combination of horizontal and vertical collaboration, resulting in hybrid collaboration, as illustrated in Fig. 16. For instance, when addressing highly complex tasks, the problem is first broken down into manageable sub-tasks, each assigned to specialized LM agents. Horizontal collaboration allows agents to perform parallel evaluations or validations, while vertical collaboration ensures that the task is refined through sequential processing stages. This computational collaboration blends both paradigms, coordinating parallel assessments across agents while sequentially integrating their outputs through defined stages, thus optimizing task execution and enhancing overall system performance. For instance, Chen et al. [111] propose a dynamic multi-agent collaboration framework that organizes agents using directed acyclic graphs to facilitate interactive reasoning. Their framework demonstrates superior performance across various network topologies and enables collaboration among thousands of agents. A key finding in [111] is the discovery of the collaborative scaling law, where solution quality improves in a logistic growth pattern as more agents are added, with collaborative emergence occurring faster than neural emergence.
② Knowledge update. Knowledge alignment is a prerequisite to update knowledge among LM agents. Zhang et al. [115] propose a fully automatic KG alignment method, using LLMs to identify and align entities in different KGs, aiming to solve the heterogeneity problem between different KGs and integrate multi-source data. (i) For explicit knowledge updates, Tandon et al. [116] pair LMs with a growing memory to train a correction model, where users identify output errors and provide general feedbacks on how to correct them. (ii) For implicit knowledge updates, continual learning ensures that AI model can continuously learn while receiving new tasks and new data without forgetting previously learned knowledge. Qin et al. [117] propose ELLE, which flexibly extends the breadth and depth of existing PLMs, allowing the model to continuously grow as new data flows in.
① Knowledge integration, which integrates and optimizes knowledge from different LM agents to form a robust and comprehensive public knowledge base for a common task. Jiang et al. [118] propose an ensemble framework named LLM-blender that directly aggregates the outputs of different models to improve prediction performance and robustness. However, it requires maintaining multiple trained LLMs and executing each LLM during inference, which can be impractical for LM agents. To address this issue, Wan et al. [79] propose an implicit knowledge fusion framework that evaluates the predictive probability distributions of multiple LLMs and uses the distribution for continuous training the target model.
@@
TABLE VI: A Summary of Knowledge Cooperation Approaches for LM Agents
@@
Online knowledge distillation LMs generate high-quality reasoning processes for models to learn. [113]
@@
Knowledge Sharing Explicit knowledge update Adjust LM via knowledge bases and feedbacks. [116]
@@
Use real-time updated knowledge sources for accurate reasoning.
[80, 78, 77]
@@
c) Knowledge retrieval. LM agents not only rely on the learned knowledge during pre-training but also can dynamically access and query external knowledge bases (e.g., databases, the Internet, and KGs) to obtain the latest information to help reasoning. RAG technology combines information retrieval and generation models by first retrieving relevant contents and then generating answers based on these contents. According to the data type, it can be divided into two categories.
@@
  • RAG based on static knowledge sources, such as Wikipedia and documents. Lewis et al. [58] demonstrate how RAG generates accurate and contextually relevant responses by retrieving relevant knowledge chunks from static sources. For retrieval enhancement, Trivedi et al. [59] propose a new multi-step QA method named IRCoT to interleave retrieval with steps in CoT, which first uses CoT to guide retrieval and then uses retrieval results to improve CoT.
  • RAG based on dynamic knowledge sources, such as news APIs. It contains three lines: exploring new knowledge, retrieving past knowledge, and self-guided retrieval. (i) For new knowledge exploration, Dai et al. [80] use RAG technology for improved safety and reliability of autonomous driving systems by utilizing real-time updated data sources, including in-vehicle sensor data, traffic information, and other driving-related dynamic data. It demonstrates RAG’s potential in handling complex environments and responding to unexpected situations. (ii) For past knowledge retrieval, Kuroki et al. [78] develop a novel vector database named coordination skill database to efficiently retrieve past memories in multi-agent scenarios to adapt to new cooperation missions. (iii) For self-guided retrieval, Wang et al. [77] propose a method called self-knowledge-guided retrieval (SKGR), which harnesses both internal and external knowledge for enhanced retrieval by allowing LLMs to adaptively call external resources when handling new problems.
@@ III-G Summary and Lessons Learned
LM agent networks bridge the human, physical, and cyber worlds, with an architecture that integrates cloud, edge, and end devices. Powered by advanced technologies including foundation models, knowledge-related technologies, interaction technologies, multi-agent collaboration, and digital twin technologies, LM agent networks enable seamless communication, sophisticated reasoning, and scalable task execution. By leveraging hierarchical frameworks, semantic-aware communication, information-centric routing, and spatiotemporal awareness, LM agent networks optimize resource allocation, enhance real-time decision-making, and improve both intra- and inter-agent communication.
@@
Figure 20: The taxonomy of security threats to LM agent networks.
@@ IV-A Authentication Threats in LM Agent Networks
Efficient and reliable authentication mechanisms is a prerequisite for safeguarding the functionality and integrity of LM agent systems. Based on the interacting entities, authentication threats to LM agents can be categorized into three types: user-agent authentication threats, agent-agent authentication threats, and agent-TPSP (third-party service provider) authentication threats, as illustrated in Fig. 21.

### v3

@@ III-E2 Distributed Competition
① MAD. The cognitive behaviors of LLMs, such as self-reflection, have proven effective in solving NLP tasks but can also lead to thought degeneration due to biases, rigidity, and limited feedback. MAD [109] introduces structured arguments and debates among LM agents, where they defend their positions and challenge each other’s strategies. By engaging in a dynamic tit-for-tat process, LM agents can correct biases, overcome resistance to change, and provide mutual feedback, leading to improved reasoning and decision-making. In MAD, diverse role prompts (i.e., personas) play a crucial role, shaping agents’ perspectives and debate strategies. Generally, as illustrated in Fig. 21, there are three primary communication strategies:
  • One-by-one debate strategy: Agents respond sequentially in a fixed order, generating replies based on the ongoing discussion. Each agent’s chat history is updated with previous responses to maintain context.
  • Simultaneous-talk debate strategy: Agents respond asynchronously within each round, mitigating the influence of speaking order on the debate’s flow.
  • Simultaneous-talk-with-summarizer debate strategy: The simultaneous-talk-with-summarizer enhances simultaneous-talk by incorporating a summarizer LLM. After each round, the summarizer condenses the discussion into a cohesive summary, replacing individual chat histories to ensure a shared and coherent context for all agents.
② Non-cooperative game. In non-cooperative games, LM agents engage in strategic decision-making where each LM agent’s goal is to seek the Nash equilibrium. Various non-cooperative games can be employed such as Nash bargaining, auctions, matching games. For instance, Xu et al. [110] design a new cached model-as-a-resource paradigm and introduce the concept of age-of-thought to efficiently deploy LLM agents. They also devise a deep Q-network-based auction mechanism to incentivize network operators.
@@ III-E4 Lessons Learned
In distributed collaboration, SL and MoE focus on model segmentation to significantly reduce computational burdens on edge devices, while FL and role-playing emphasize model integration to enhance cooperation among distributed agents. Within the overall cloud-edge collaboration for LM agent systems, key challenges emerge in the joint consideration of LM segmentation (e.g., SL and MoE), LM placement, LM replacement, LM routing, and knowledge caching, alongside the inherent heterogeneity and dynamics of edge environments. Effective deployment and operation of LM agents across cloud and edge infrastructures require a delicate balance between LM segmentation and LM integration strategies, with an emphasis on optimizing resource allocation and minimizing service latency. Additionally, solutions should be adaptable to the constantly changing network conditions and hardware limitations inherent to edge devices. Future research includes the seamless integration of these techniques, taking into account both static and dynamic aspects of edge networks, to offer more efficient and scalable solutions for large-scale applications.
@@ III-F Cooperation Paradigms of Connected LM Agents
As illustrated in Fig. 22 and Table VII, the detailed cooperation paradigms of LM agents under cloud-edge-end architecture involve three perspectives: data cooperation, computation cooperation, and knowledge cooperation.
  • Data Cooperation: Within a common task, LM agents continuously exchange and fuse their individual data (e.g., task-oriented sensory data) to ensure a comprehensive and up-to-date understanding of their task environment, thereby enhancing collective intelligence and enabling coordinated actions.
@@
a) Multimodal Data Cooperation emphasizes the fusion of data from various modalities, such as text, images, audio, and video, to offer a comprehensive understanding of the environment. This cooperation allows LM agents to process and interpret information from multiple sources, leading to more accurate and robust decision-making. (i) By combining data from different modalities, it helps to create a unified representation which harnesses the strengths of each type of data. For instance, Gross et al. [113] discuss the use of multimodal data to model communication in artificial social agents, emphasizing the importance of verbal and nonverbal cues for natural human-robot interaction. (ii) By enabling LM agents to retrieve relevant information across different modalities, it enhances their ability to respond to complex queries and scenarios. For instance, by considering the intra-modality similarities in multi-modal video representations, Zolfaghari et al. [114] introduce the contrastive loss in contrastive learning process for enhanced cross-modal embedding, whose effectiveness is validated using LSMDC and YouCook2 datasets for video-text retrieval and video captioning tasks.
@@
a) Horizontal Collaboration. It refers to decompose complex tasks into manageable sub-tasks, with multiple LM agents independently completing their assigned tasks in parallel. The results are then summarized and integrated to generate the final outcome. This approach enables the collaborative system to scale horizontally by adding more LM agents, allowing for more complex and dynamic tasks to be handled. The parallel processing and multi-angle perspectives contribute to increased robustness and reduced errors and biases that may arise from relying on a single agent. As shown in Fig. 23(a), each LM agent independently analyzes different aspects of stock performance such as trend, news, and industry to assess its investment potential. Horizontal collaboration has been successfully implemented in various LM agent systems. For instance, ProAgent [86] employs a framework where LM agents collaborate as teammates, analyzing each other’s intentions and updating their beliefs based on observed behaviors of their peers. This enhances collective decision-making by allowing agents to adapt dynamically in real time. Similarly, DyLAN [111] assembles a team of strategic agents that communicate through task-specific queries, allowing multiple rounds of interaction to enhance both efficiency and overall performance. However, the independent analysis of each LM agent may lead to inconsistent outputs, complicating the aggregation of the final conclusions. Therefore, effective mechanisms are needed to address disagreements and ensure that the agents’ contributions are complementary, which may increase the complexity of the coordination process.
@@
b) Vertical Collaboration. It involves decomposing complex tasks into multiple stages, with different LM agents handling each stage sequentially. After completing their respective tasks, each agent passes the result to the next agent, until the entire task is successfully completed. As illustrated in Fig. 23(b), in medical image analysis, the first agent might extract basic visual features, the second agent could analyze these features in greater depth, and the final agent produces a diagnosis. This sequential approach enables a step-by-step refinement, allowing complex problems to be broken down and tackled more efficiently by leveraging specialized agents at each stage. For instance, Jiang et al. [18] present CommLLM, a multi-agent system for natural language-based communication tasks. It comprises three key components: i) multi-agent data retrieval, utilizing condensate and inference agents to refine 6G communication knowledge; ii) multi-agent collaborative planning, employing various planning agents to generate solutions from multiple viewpoints; and iii) multi-agent reflection, which evaluates solutions and suggests improvements through reflexion and refinement agents. However, a limitation of vertical collaboration is that higher-level LM agents rely on the accuracy of lower-level outputs, making the process vulnerable to error propagation. Mistakes made early in the sequence can compromise the final outcome, highlighting the need for high precision at each stage to maintain the overall system’s reliability.
@@
TABLE VII: A Summary of Data, Computation, and Knowledge Cooperation Approaches for LM Agents
@@
Data Multimodal Data Cooperation
Modality alignment
& merging
Use multimodal data to model communication in social AI agents. [113]
Multimodal retrieval Utilize contrastive learning for cross-modal embedding. [114]
Spatio-temporal Data Cooperation Collaborative perception Combine temporal semantics & spatial agent data for multi-source fusion. [88]
Collaborative prediction Combine temporal & spatial graph convolutions for traffic prediction. [115]
Computation Horizontal Collaboration Role-playing Agents cooperate with diverse roles: Controller, Planner, & Verificator. [86]
Dynamic agent team Agents dynamically form a team via agent selection and evaluation. [111]
Vertical Collaboration CoT reasoning Agents cooperatively reason complex task via CoT. [44]
Chained cooperation Multi-agent cooperation through data retrieval, planning, and reflection. [18]
Hybrid Collaboration ToT &GoT reasoning Agents cooperatively reason complex task via ToT and GoT. [45, 46]
DAG reasoning Organize agents as DAGs for interactive reasoning. [116]
Knowledge Knowledge Sharing Online knowledge distillation LMs generate high-quality reasoning processes for models to learn. [54]
@@
Explicit knowledge update Adjust LM via knowledge bases and feedbacks. [118]
@@
Dynamic knowledge base Use real-time updated knowledge sources for accurate reasoning. [82, 80, 79]
c) Hybrid collaboration. In practical LM agent environments, real-world applications often require a combination of horizontal and vertical collaboration, resulting in hybrid collaboration, as illustrated in Fig. 20. For instance, when addressing highly complex tasks, the problem is first broken down into manageable sub-tasks, each assigned to specialized LM agents. Horizontal collaboration allows agents to perform parallel evaluations or validations, while vertical collaboration ensures that the task is refined through sequential processing stages. This computational collaboration blends both paradigms, coordinating parallel assessments across agents while sequentially integrating their outputs through defined stages, thus optimizing task execution and enhancing overall system performance. For instance, Chen et al. [116] propose a dynamic multi-agent collaboration framework that organizes agents using directed acyclic graphs (DAGs) to facilitate interactive reasoning. Their framework demonstrates superior performance across various network topologies and enables collaboration among thousands of agents. A key finding in [116] is the discovery of the collaborative scaling law, where solution quality improves in a logistic growth pattern as more agents are added, with collaborative emergence occurring faster than neural emergence.
② Knowledge update. Knowledge alignment is a prerequisite to update knowledge among LM agents. Zhang et al. [124] propose a fully automatic KG alignment method, using LLMs to identify and align entities in different KGs, aiming to solve the heterogeneity problem between different KGs and integrate multi-source data. (i) For explicit knowledge updates, Tandon et al. [118] pair LMs with a growing memory to train a correction model, where users identify output errors and provide general feedbacks on how to correct them. (ii) For implicit knowledge updates, continual learning ensures that AI models can continuously learn while receiving new tasks and new data without forgetting previously learned knowledge. Qin et al. [119] propose ELLE, which flexibly extends the breadth and depth of existing PLMs, allowing the model to continuously grow as new data flows in.
① Knowledge integration, which integrates and optimizes knowledge from different LM agents to form a robust and comprehensive public knowledge base for a common task. Jiang et al. [120] propose an ensemble framework named LLM-blender that directly aggregates the outputs of multiple models for enhanced prediction accuracy and robustness. However, it necessitates maintaining several trained LLMs and running each LLM during inference, making it less practical for LM agents. To address this issue, Wan et al. [81] propose an implicit knowledge fusion framework that evaluates the predictive probability distributions of multiple LLMs and uses the distribution for continuous training the target model.
@@
c) Knowledge retrieval. LM agents not only rely on the learned knowledge during pre-training but also can dynamically access and query external knowledge bases (e.g., databases, the Internet, and KGs) to obtain the latest information to help reasoning. RAG technology combines information retrieval and generation models by first retrieving relevant contents and then generating answers based on these contents. Based on the data type, it can be divided into two categories.
@@
① RAG based on static knowledge sources, such as Wikipedia and documents. Lewis et al. [53] demonstrate how RAG generates accurate and contextually relevant responses by retrieving relevant knowledge chunks from static sources. For retrieval enhancement, Trivedi et al. [55] propose a new multi-step QA method named IRCoT to interleave retrieval with steps in CoT, which first utilizes CoT to guide retrieval and then uses retrieval outcomes to enhance CoT.
② RAG based on dynamic knowledge sources, such as news APIs. It contains three lines: exploring new knowledge, retrieving past knowledge, and self-guided retrieval.
  • For new knowledge exploration, Dai et al. [82] use RAG technology for improved safety and reliability of autonomous driving systems by utilizing real-time updated data sources, including in-vehicle sensor data, traffic information, and other driving-related dynamic data. It demonstrates RAG’s potential in handling complex environments and responding to unexpected situations.
  • For past knowledge retrieval, Kuroki et al. [80] develop a novel vector database named coordination skill database to efficiently retrieve past memories in multi-agent scenarios to adapt to new cooperation missions.
  • For self-guided retrieval, Wang et al. [79] propose a method called self-knowledge-guided retrieval (SKGR), which harnesses both internal and external knowledge for enhanced retrieval by allowing LLMs to adaptively call external resources when handling new problems.
Figure 24: The taxonomy of security threats to LM agent networks.
@@ III-G Summary and Lessons Learned
With the rapid advancement of LMs, the future will witness a proliferation of personal assistant agents and embodied agents such as robots. LM agent networks bridge the human, physical, and cyber worlds, with an architecture that integrates cloud, edge, and end devices. Powered by advanced technologies including foundation models, knowledge-related technologies, interaction technologies, multi-agent collaboration, and digital twin technologies, LM agent networks enable seamless communication, sophisticated reasoning, and scalable task execution. By leveraging hierarchical frameworks, semantic-aware communication, information-centric routing, and spatiotemporal awareness, LM agent networks optimize resource allocation, enhance real-time decision-making, and improve both intra- and inter-agent communication.
@@
  • LM agent networks leverage distributed cloud-edge-end computing resources, enabling software/hardware agents at the edge and terminal levels to collaborate efficiently on complex tasks. This decentralized approach challenges the traditional trend of building ever-larger models by promoting modular, task-specific agents that can dynamically interact and share capabilities. Such a paradigm not only enhances computational efficiency and scalability but also improves adaptability across diverse application scenarios, paving the way for more flexible and resource-efficient AI ecosystems. However, significant challenges remain in designing agent capability notification, dynamic discovery, communication protocols, task orchestration, and agent routing across diverse scenarios, while also ensuring robust security and privacy protection.
@@ IV-A Authentication Threats in LM Agent Networks
Efficient and reliable authentication mechanisms are a prerequisite for safeguarding the functionality and integrity of LM agent systems. Based on the interacting entities, authentication threats to LM agents can be categorized into three types: user-agent authentication threats, agent-agent authentication threats, and agent-TPSP (third-party service provider) authentication threats, as illustrated in Fig. 25.